### v2

@@ Abstract
Compositional generalization (the ability to respond correctly to novel combinations of familiar components) is thought to be a cornerstone of intelligent behavior. Compositionally structured (e.g. disentangled) representations are essential for this; however, the conditions under which they yield compositional generalization remain unclear. To address this gap, we present a general theory of compositional generalization in kernel models with fixed representations, a tractable framework for characterizing the impact of dataset statistics on generalization. We find that kernel models are constrained to adding up values assigned to each combination of components seen during training (“conjunction-wise additivity”). This imposes fundamental restrictions on the set of tasks these models can learn, in particular preventing them from transitively generalizing equivalence relations. Even for compositional tasks that kernel models can in principle learn, we identify novel failure modes in compositional generalization that arise from biases in the training data and affect important compositional building blocks such as symbolic addition and context dependence (memorization leak and shortcut bias). Finally, we empirically validate our theory, showing that it captures the behavior of deep neural networks (convolutional networks, residual networks, and Vision Transformers) trained on a set of compositional tasks with similarly structured data. Ultimately, this work provides a theoretical perspective on how statistical structure in the training data can affect compositional generalization, with implications for how to identify and remedy failure modes in deep learning models.
@@ 1 Introduction
Humans’ understanding of the world is inherently compositional: once familiar with the concepts “pink” and “elephant,” we can immediately imagine a pink elephant. Stitching together concepts in this way lets humans generalize far beyond our prior experience, allowing us to cope with unfamiliar situations and imagine things that do not yet exist (Lake et al., 2017; Frankland & Greene, 2020). Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem (Fodor & Pylyshyn, 1988; Battaglia et al., 2018; Lake & Baroni, 2018). While a wide range of studies have investigated the conditions under which compositionally structured (i.e. “disentangled”) representations can be learned (Hinton et al., 2011; Higgins et al., 2017; Träuble et al., 2021; Whittington et al., 2023; Ren et al., 2023), it remains unclear when learning these representations is actually useful to a downstream neural network. Some work suggests that disentangled representations improve compositional generalization (Esmaeili et al., 2019; van Steenkiste et al., 2019). However, others challenge this view (Locatello et al., 2019; Schott et al., 2022). In general, it remains unclear why insights from some compositional tasks fail to generalize to others, and a systematic understanding of the relationships among compositional tasks remains elusive (Hupkes et al., 2020).
@@
To make progress on this question, we focus on standard statistical learning, a fundamental basis for compositional generalization that affects almost any machine learning model. Generalization in statistical learning depends on the similarities between different inputs. Notably, in compositionally structured representations, inputs that have components in common (say, a red circle and a blue circle) are more similar to each other than inputs that do not (say, a red circle and a blue square). As a result, the compositional structure of a task is reflected in its dataset statistics, influencing the generalization behavior of the model.
@@
To understand when and how statistical learning leads to successful compositional behavior, we developed a theory for the behavior of kernel models on compositional tasks. Kernel models are an important class of statistical models that are able to provide a simplified approximation to neural networks while maintaining analytical tractability (Jäkel et al., 2008, 2009). For example, kernel models accurately describe the behavior of learning and fine-tuning in neural networks under certain conditions (Jacot et al., 2018; Malladi et al., 2023). More broadly, they provide a tractable framework for characterizing the impact of dataset statistics on generalization (Canatar et al., 2021a, b).
@@
Despite their broad relevance and relative simplicity, it has remained unclear how even kernel models generalize on compositional tasks (though see Abbe et al., 2023; Lippl et al., 2024, see Section 2). To address this gap, we present a theory of compositional generalization in kernel models. We define a general family of compositional tasks with no constraints on the input-output mapping (ensuring broad applicability of our theory), theoretically characterize the compositional behavior of kernel models trained on such tasks, and go beyond kernel models to empirically validate our theory in several relevant deep neural network architectures. Our specific contributions are as follows:
@@
  • In Section 4, we show that kernel models with disentangled inputs are constrained to summing up values implicitly assigned to each component or combination of components seen during training. We call such computations “conjunction-wise additive.”
@@
  • In Section 5, we then characterize how representational geometry determines whether kernel models will generalize successfully on conjunction-wise additive compositional tasks, highlighting two important failure modes (memorization leak and shortcut bias). This highlights that disentangled representations are not sufficient for downstream compositional generalization (even on conjunction-wise additive tasks) and explains why.
@@
Overall, we take a step towards a general theory of compositional generalization by characterizing the impact of dataset statistics on compositional generalization for a broad range of tasks. Our theory systematically clarifies the compositional generalization behavior of deep neural networks and lays a foundation for the design of new learning mechanisms that overcome their limitations.
@@ 2 Related work
Compositional generalization. Compositionality is an important theme in both human and machine reasoning problems, including visual reasoning (Lake et al., 2015; Johnson et al., 2017; Schwartenbeck et al., 2023), language production (Hupkes et al., 2020), and rule learning (Ito et al., 2022; Abdool et al., 2023). Recent breakthroughs have led to massive improvements in models’ compositional capacities, but in some cases, these models still fail spectacularly (Srivastava et al., 2023; Lewis et al., 2023; West, 2023). Attempts to improve compositional generalization often leverage meta-learning (Mitchell et al., 2021; Wu et al., 2023; Lake & Baroni, 2023) or modular architectures (Andreas et al., 2017), in the hopes that different modules will specialize for different components (Schug et al., 2024). Constraining a network to be linear (or additive) can guarantee modular specialization and compositional generalization (Lachapelle et al., 2023; Wiedemer et al., 2023a, b). However, these networks are extremely limited in the tasks they can learn (as we will show below) and end-to-end training of modular architectures without such constraints often does not result in the desired specialization (Bahdanau et al., 2019; Mittal et al., 2022; Jarvis et al., 2023).
Conjunctions. Henceforth, we will call a combination of components a conjunction. Conjunctive codes have a long history in neuroscience (Alvarado & Rudy, 1992; Baker et al., 2002), and are theoretically linked to the formation of highly specific episodic memories.
@@
Kernel regime. When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(z)=w^{T}\phi(z)$, it can be shown that the resulting behavior of that model can be described in terms of the kernel $K(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$ induced by the representation $\phi(z)$ (Schölkopf, 2000). We will refer to such models as kernel models. Notably, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, a kernel model can be written in its “dual form” as
$f_{a}(z)=\sum_{i=1}^{n}a_{i}K(z,z_{i}),\mbox{ for dual coefficients}a\in\mathbb{R}^{n}.$ (1)
Prior work has shown that neural networks with large initial weights or wide architectures are well approximated by gradient descent on a model with a fixed representation (a kernel model) (Jacot et al., 2018). This is called the “kernel regime,” in contrast to the feature-learning regime, which is brought forth, for example, by small initial weights or small width (Chizat et al., 2019).
@@
Norm minimization. Gradient descent, ridge regression, and neural networks in the kernel regime learn the readout weights that describe the training data with minimal $\ell_{2}$-norm (Soudry et al., 2018; Gunasekar et al., 2018; Ji et al., 2020). Norm minimization is a standard theoretical framework for analyzing how representational geometry and dataset statistics influence generalization (Canatar et al., 2021b, 2023) and we here apply it to the compositional task setting. This is similar to Lippl et al. (2024) who characterize model behavior on a specific compositional task (transitive ordering) and Abbe et al. (2023) who characterize the inductive bias of norm minimization for inputs with binary components in the limit of infinite components. Compared to these prior works, we analyze a broader range of compositional tasks and derive exact constraints for finite numbers of components.
@@
Shortcut learning. Shortcut learning refers to models exploiting spurious correlations between certain features of the input data and the target (Shah et al., 2020; Nagarajan et al., 2021). This substantially impacts their performance out of distribution (where those correlations may not hold) (Geirhos et al., 2020). We theoretically analyze how shortcut biases impact compositional generalization.
@@ 3.1 Task space
We assume that our model extracts from its input a set of disentangled components $(z_{c})_{c=1}^{C}$, e.g. color, shape, and background (Fig. 1a). For instance, the model could have been pre-trained to extract these components from an image or an input sentence. Each component $z_{c}\in Z_{c}\subset\mathbb{R}^{d}$ is drawn from a discrete set of possible components represented by $d$-dimensional real vectors and because we consider categorical components, we assume that all components have unit magnitude and are orthogonal to each other (these assumptions can be relaxed, Section A.1). We then assume that the model represents these components by adding them together: $z:=\sum_{c=1}^{C}z_{c}$. A simple example of this is given by a concatenation of one-hot representations of all components (Fig. 1a). The target, $y\in\mathbb{R}$, is given by an arbitrary function of $z$, ensuring that our framework is agnostic to the underlying compositional structure. After training models on certain combinations of components $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$.
@@
Our theory applies to any task in this family; that is, any task with a fixed number of categorical input components and a scalar output, regardless of the input-output mapping. Below we describe three example tasks within this general framework that represent important building blocks for compositional reasoning in machine learning and cognitive science. We will focus on these tasks in the main text, to illustrate our general theory and analyze its consequences in more detail. In Appendix D, we describe additional tasks captured by our theory.
@@
Symbolic addition. Many tasks involve inferring a magnitude associated with different underlying components (e.g. adding handwritten digits) (Lorenzi et al., 2021; Sheahan et al., 2021). We consider two components $(z_{1},z_{2})$ with unobserved assigned values $v_{1}(z_{1})$ and $v_{2}(z_{2})$. The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$. After sufficient exposition to individual items, a model with an additive structure can generalize to novel combinations of items. In particular, we consider nine input elements $[-4],[-3],\dotsc,[4]$ with associated values $-4,-3,\dotsc,4$ and training sets containing all pairs where at least one component is equal to a certain subset of values (given by the rows and columns in the training set in Fig. 1b). We vary the size of this subset (i.e. the number of rows/columns) as well as whether the task requires only interpolation or also extrapolation.
@@
Context dependence. The relevance of different stimuli often depends on our current context. Taking this into account is a crucial aspect of cognition in humans and animals (Bouton et al., 1999; Taylor & Ivry, 2013; Parker & Hollister, 2014; Ito et al., 2022). We therefore consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$. The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response. Both features have six possible values which are split up in two categories. If the model has learned this context dependence, it should be able to generalize to novel feature combinations. We evaluate on the subset of data for which $z_{f1}$ indicates Cat. 2 and $z_{f2}$ indicates Cat. 1. In the most extreme generalization test (CD-3), we leave out the entire subset; in easier versions, we leave out combinations of two or one of those features (CD-2 and CD-1) (Fig. 1c).
@@
Transitive equivalence. Relational reasoning is an important example of compositional generalization and often involves extending learned relations to new item combinations (Halford et al., 2010; Battaglia et al., 2018). Given an unobserved (and arbitrary) equivalence relation, the task is to determine whether two presented items $(z_{1},z_{2})$ are equivalent. The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig. 1d). This is an important instance of relational cognition (often studied as “associative inference” in cognitive science (Schlichting & Preston, 2015; Spalding et al., 2018)). Although prior work has found that kernel models often successfully generalize on a transitive ordered relation ($A>B$ and $B>C$ imply $A>C$, Lippl et al., 2024), the behavior of kernel models on equivalence relations has remained unclear.
@@ 3.2 Models
Kernel models. Our theory characterizes kernel models with a disentangled input $z\in\mathbb{R}^{d}$ that apply a transform $\phi(z)\in\mathbb{R}^{h}$ and learn a linear readout, $f_{w}(x):=w^{T}\phi(z)$, using gradient descent or ridge regression (Fig. 1a). For $\phi$, we consider a neural network with random weights (in the infinite-width limit; see Section A.3). Our model setup includes as a special case the random feature model studied by Abbe et al. (2023) and further captures training via backpropagation in the kernel regime.
Deep neural networks. Additionally, we consider a broad range of deep neural network architectures: convolutional networks, residual networks (He et al., 2016), Vision Transformers (ViTs; Dosovitskiy et al., 2020)), and ReLU networks; all trained through backpropagation. The code required to reproduce all experiments can be found at https://github.com/sflippl/compositional-generalization.
Figure 1: a, Theoretical setup: a disentangled representation of the input, followed by a nonlinear transform $\phi$ and a learned linear readout $w$. b, Example training sets for symbolic addition. The grid represents nine components with associated values $-4,-3,\dotsc,4$. The training set consists of certain rows and columns of this grid. c, Context dependence. In context 1, feat. 1 determines the category; in context 2, feat. 2 determines the category. The training sets leave out different subsets of the lower right orthant. d, Transitive equivalence: six items are split up into two arbitrary equivalence classes (e.g. A,B,C and D,E,F) and generalization requires transitive inference over equivalence classes.

### v3

@@ Abstract
Compositional generalization (the ability to respond correctly to novel combinations of familiar components) is thought to be a cornerstone of intelligent behavior. Compositionally structured (e.g. disentangled) representations support this ability; however, the conditions under which they are sufficient for the emergence of compositional generalization remain unclear. To address this gap, we present a theory of compositional generalization in kernel models with fixed, compositionally structured representations. This provides a tractable framework for characterizing the impact of training data statistics on generalization. We find that these models are limited to functions that assign values to each combination of components seen during training, and then sum up these values (“conjunction-wise additivity”). This imposes fundamental restrictions on the set of tasks compositionally structured kernel models can learn, in particular preventing them from transitively generalizing equivalence relations. Even for compositional tasks that they can learn in principle, we identify novel failure modes in compositional generalization (memorization leak and shortcut bias) that arise from biases in the training data. Finally, we empirically validate our theory, showing that it captures the behavior of deep neural networks (convolutional networks, residual networks, and Vision Transformers) trained on a set of compositional tasks with similarly structured data. Ultimately, this work examines how statistical structure in the training data can affect compositional generalization, with implications for how to identify and remedy failure modes in deep learning models.
@@ 1 Introduction
Humans’ understanding of the world is inherently compositional: once familiar with the concepts “pink” and “elephant,” we can immediately imagine a pink elephant. Stitching together concepts in this way lets humans generalize far beyond our prior experience, allowing us to cope with unfamiliar situations and imagine things that do not yet exist (Fig. 1a) (Lake et al., 2017; Frankland & Greene, 2020). Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem (Fodor & Pylyshyn, 1988; Battaglia et al., 2018; Lake & Baroni, 2018; Lepori et al., 2023). While a wide range of studies have investigated the conditions under which compositionally structured (i.e. “disentangled”) representations can be learned (Hinton et al., 2011; Higgins et al., 2017; Träuble et al., 2021; Whittington et al., 2023; Ren et al., 2023), it remains unclear when learning these representations is actually useful to a downstream neural network. Some work suggests that disentangled representations improve compositional generalization (Esmaeili et al., 2019; van Steenkiste et al., 2019). Other studies challenge this view (Locatello et al., 2019; Schott et al., 2022). In general, a systematic understanding of the relationships among compositional tasks remains elusive (Hupkes et al., 2020).
Figure 1: Overview of main theoretical findings (Section 4). a, We consider inputs with several categorical components and study compositional generalization to novel combinations of components. b, We assume compositionally structured representations for which trials with the same number of overlaps have identical similarities. c, We find that a random weights neural network conserves compositional structure: if its input is compositionally structured then so is its output. d, We find that compositionally structured kernel models are constrained to adding up values for each conjunction seen during training (“conjunction-wise additivity”).
@@
To make progress on this question, we focus on standard statistical learning, a fundamental basis for generalization that affects almost any machine learning model. Generalization in statistical learning depends on the similarities between different inputs. In compositionally structured representations, inputs that have components in common (say, a red circle and a blue circle) are more similar to each other than inputs that do not (say, a red circle and a blue square). As a result, the compositional structure of a task is reflected in its dataset statistics, influencing the model’s generalization behavior.
@@
To understand when and how statistical learning leads to successful compositional behavior, we developed a theory for the behavior of kernel models on compositional tasks. Kernel models are an important class of statistical models that are able to provide a simplified approximation to neural networks while maintaining analytical tractability (Jäkel et al., 2008; 2009). For example, kernel models accurately describe the behavior of learning and fine-tuning in neural networks under certain conditions (Jacot et al., 2018; Malladi et al., 2023). More broadly, they provide a tractable framework for characterizing the impact of dataset statistics on generalization (Canatar et al., 2021a; b).
@@
Despite their broad relevance and relative simplicity, it has remained unclear how even kernel models generalize on compositional tasks (though see Abbe et al., 2023; Lippl et al., 2024, see Section 2). To address this gap, we present a theory of compositional generalization in kernel models. We define a general class of “compositionally structured representations” (Fig. 1b) and a general family of compositional tasks with no constraints on the input-output mapping (ensuring broad applicability of our theory). We then theoretically characterize the compositional behavior of compositionally structured kernel models, and go beyond kernel models to empirically validate our theory in several relevant deep neural network architectures. Our specific contributions are as follows:
@@
  • In Section 4, we show that compositionally structured kernel models are constrained to summing up values implicitly assigned to each component or combination of components seen during training (Fig. 1). We call such computations “conjunction-wise additive.”
@@
  • In Section 5, we then characterize how representational geometry determines whether kernel models will generalize successfully on conjunction-wise additive compositional tasks, highlighting two important failure modes (memorization leak and shortcut bias). This demonstrates that disentangled representations are not sufficient for downstream compositional generalization (even on conjunction-wise additive tasks) and explains why.
@@
Overall, we take a step towards a general theory of compositional generalization. Our theory systematically clarifies the compositional generalization behavior of deep networks and lays a foundation for the design of new learning mechanisms that overcome their limitations.
@@ 2 Related work
Compositional generalization. Compositionality is an important theme in both human and machine cognition, including visual reasoning (Lake et al., 2015; Johnson et al., 2017; Schwartenbeck et al., 2023), language production (Hupkes et al., 2020), and rule learning (Ito et al., 2022; Abdool et al., 2023). While recent breakthroughs have led to massive improvements in models’ compositional capacities, these models can still fail spectacularly (Srivastava et al., 2023; Lewis et al., 2023; West, 2023). Attempts to improve compositional generalization often leverage meta-learning (Mitchell et al., 2021; Wu et al., 2023; Lake & Baroni, 2023) or modular architectures (Andreas et al., 2017). Constraining a network’s compositional function (e.g. to adding up a value for each component) can guarantee modular specialization and compositional generalization (Lachapelle et al., 2023; Wiedemer et al., 2023a; b; Schug et al., 2024). However, these networks are extremely limited in the tasks they can learn (as we will show below) and end-to-end training of modular architectures without such constraints often does not result in the desired specialization (Bahdanau et al., 2019; Mittal et al., 2022; Jarvis et al., 2023).
@@
Kernel regime. When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(x)=w^{T}\phi(x)$, the behavior of that model can be described in terms of the kernel $K_{\phi}(x,x^{\prime})=\phi(x)^{T}\phi(x^{\prime})$ induced by $\phi(x)$ (Schölkopf, 2000). Notably, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, such “kernel models” can be written in dual form as $f_{a}(x)=\sum_{i=1}^{n}a_{i}K_{\phi}(x,x_{i})$, for inferred dual coefficients $a\in\mathbb{R}^{n}$. Prior work has shown that neural networks with large initial weights or wide architectures are well approximated by a kernel model (“kernel regime”) (Jacot et al., 2018). In contrast, the “feature-learning regime” (which can be brought forth, for example, by small initial weights or small width) yields more substantial changes in neural networks’ internal representations (Chizat et al., 2019).
@@
Norm minimization. Gradient descent, ridge regression, and neural networks in the kernel regime learn the readout weights that describe the training data with minimal $\ell_{2}$-norm (Soudry et al., 2018; Gunasekar et al., 2018; Ji et al., 2020). Norm minimization is a standard theoretical framework for analyzing how representational geometry and dataset statistics influence generalization (Canatar et al., 2021a; b; 2023) and we here apply it to the compositional task setting. This is similar to Lippl et al. (2024) who characterize model behavior on a specific compositional task (transitive ordering) and Abbe et al. (2023) who characterize the inductive bias of norm minimization for inputs with binary components in the limit of infinite components. Compared to these prior works, we analyze a broader range of compositional tasks and derive exact constraints for finite numbers of components.
@@
Shortcut learning. Shortcut learning refers to models exploiting spurious correlations between certain features of the input data and the target (Shah et al., 2020; Nagarajan et al., 2021). This substantially impacts their performance out of distribution, where those correlations may not hold (Geirhos et al., 2020). We theoretically analyze how shortcut biases impact compositional generalization.
@@ 3.1 Task space
We consider an input $x$ representing a set of underlying components $z=(z_{c})_{c=1}^{C}$, where each component $z_{c}\in Z_{c}$ is drawn from a discrete, finite set of possible components. For example, $x$ could be a “multi-hot” concatenation of one-hot representations of all components (Fig. 1c). We make a specific assumption about how the trial-by-trial similarity of input representations $x$ relates to the underlying components $z$ (Fig. 1b):
Definition 3.1.
A representation $x$ is “compositionally structured” iff its kernel $K(x,x^{\prime})=x^{T}x^{\prime}$ only depends on the number of components that are identical between $z$ and $z^{\prime}$, where $z$ and $z^{\prime}$ are the components represented by $x$ and $x^{\prime}$ respectively.
In particular, the multi-hot representation described above is compositionally structured, but we will see below that this concept captures a richer set of representations as well. The target, $y\in\mathbb{R}$, is given by an arbitrary function of $z$, ensuring that our framework is agnostic to the underlying compositional structure. After training models on certain combinations of components $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$.
@@
Our theory characterizes constraints on model behavior for arbitrary tasks within this family, regardless of the ground-truth input-output mapping. Below we describe three example tasks that represent important building blocks for compositional reasoning in machine learning and cognitive science. We will focus on these tasks in the main text, to illustrate our general theory and analyze its consequences in more detail. In Appendix E, we describe additional tasks captured by our theory (Fig. 2d).
Figure 2: The compositional task space. a, Example training sets for symbolic addition. The grid represents pairs of components with associated values $-4,\dotsc,4$. The training set consists of certain rows and columns of this grid. b, Context dependence. In context 1, feat. 1 determines the category; in context 2, feat. 2 determines the category. The training sets leave out subsets of the lower right orthant. c, Transitive equivalence: six items are split up into two equivalence classes (e.g. A,B,C and D,E,F) and generalization requires transitive inference over equivalence classes (e.g. $A=B$ and $B=C$ implies $A=C$). d, Our theory partitions the task space into conjunction-wise additive (which can be solved by kernel models) and non-additive tasks (which cannot).
@@
Symbolic addition. Many tasks involve inferring a magnitude associated with different underlying components (e.g. adding handwritten digits) (Lorenzi et al., 2021; Sheahan et al., 2021). We consider two components $(z_{1},z_{2})$ with hidden values $v_{1}(z_{1})$ and $v_{2}(z_{2})$. The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$. After sufficient exposure to individual items in different combinations, an additive model could in principle generalize to novel combinations. We consider training sets consisting of entire rows and columns of the compositional dataset (Fig. 2a), varying the number of rows and columns as well as whether generalization involves interpolation or also extrapolation.
@@
Context dependence. The relevance of different stimuli often depends crucially on our current context (Bouton et al., 1999; Taylor & Ivry, 2013; Parker & Hollister, 2014; Ito et al., 2022). We consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$. The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response. Features have six possible values which are divided into two classes. If the model has learned this context dependence, it should generalize to novel feature combinations. We evaluate on the subset of data for which $z_{f1}$ indicates Cat. 2 and $z_{f2}$ indicates Cat. 1. In the most extreme generalization test (CD-3), we leave out the entire subset; in easier versions, we leave out combinations of one or two features (CD-1, CD-2) (Fig. 2b).
@@
Transitive equivalence. Relational reasoning involves extending learned relations to new item combinations (Halford et al., 2010; Battaglia et al., 2018). Given an unobserved (and arbitrary) equivalence relation, the task here is to determine whether two presented items $(z_{1},z_{2})$ are equivalent. The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig. 2c). This is an important instance of relational cognition (often studied as “associative inference” in cognitive science (Schlichting & Preston, 2015; Spalding et al., 2018)). Although prior work has found that kernel models often successfully generalize on a transitive ordered relation ($A>B$ and $B>C$ imply $A>C$, Lippl et al., 2024), the behavior of kernel models on equivalence relations has remained unclear.
@@ 3.2 Models
Our theory characterizes kernel models with a compositionally structured input $x\in\mathbb{R}^{d}$ that apply a transform $\phi(x)\in\mathbb{R}^{h}$ and learn a linear readout, $f_{w}(x):=w^{T}\phi(x)$, using gradient descent with initial weights $w_{0}=0$. For $\phi$, we consider a neural network with random weights (in the infinite-width limit; see Section A.3). Our setup captures, for example, the random feature model studied by Abbe et al. (2023) and training via backpropagation in the kernel regime.